ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 253, 2023

Abadias, L., J. E. Galé and C. Lizama Poisson equation and discrete one-sided Hilbert transform for (C, α)-bounded operators

917

Aslanyan, V., S. Eterović and J. Kirby A closure operator respecting the modular j-function

321

Avitabile, M., A. Caranti, N. Gavioli, V. Monti, M. F. Newman and E. A. O’Brien Thin subalgebras of Lie algebras of maximal class

101

Besser, A. and S. L. Zerbes Vologodsky integration on curves with semi-stable reduction

761

Brändén, P., J. Leake and I. Pak Lower bounds for contingency tables via Lorentzian polynomials

43

Brendel, J., J. Kim and J. Moon On the topology of real Lagrangians in toric symplectic manifolds

113

Brešar, M. and P. Šemrl The Waring problem for matrix algebras  381

Caranti, A. See M. Avitabile

Christensen, J. D. and E. Wu Exterior bundles in diffeology  673

Ciaci, S., J. Langemets and A. Lissitsin A characterization of Banach spaces containing ℓ_{1}(κ) via ball-covering properties

359

Coskun, I. and E. Riedl Algebraic hyperbolicity of very general surfaces

787

Das, S. On the Inertia Conjecture and its generalizations  157

Detomi, E., M. Morigi and P. Shumyatsky Bounding the order of a verbal subgroup in a residually finite group

771

Dymond, M. and V. Kaluža Highly irregular separated nets  501

Eskenazis, A. and P. Ivanisvili Sharp growth of the Ornstein– Uhlenbeck operator on Gaussian tail spaces

469

Eterović, S. See V. Aslanyan Galé, J. E. See L. Abadias Gavioli, N. See M. Avitabile Geschke, S., J. Kurkofka, R. Melcher and M. Pitz Halin’s end

degree conjecture

617

González-Riquelme, C. A. On the continuity of maximal operators of convolution type at the derivative level

745

Hübner, K. Tame and strongly étale cohomology of curves  1

Hambardzumyan, L., H. Hatami and P. Hatami Dimension-free bounds and structural results in communication complexity

555

Hatami, H. See L. Hambardzumyan Hatami, P. See L. Hambardzumyan Hayut, Y. and S. Müller Perfect subtree property for weakly

compact cardinals

865

Hazrat, R. and Z. Mesyan Graded semigroups  249

Ivanisvili, P. See A. Eskenazis

Ivanov, S. O. On the cokernel of the Baumslag rationalization  407

Janzer, O. Rainbow Turán number of even cycles, repeated patterns and blow-ups of cycles

813

Kaluža, V. See M. Dymond

Karoubi, M. Real K-theories  91

Kim, J. See J. Brendel Kirby, J. See V. Aslanyan Kurkofka, J. See S. Geschke Langemets, J. See S. Ciaci Leake, J. See P. Brändén

Li, P., J. Pan and J. Wu On modular cohomotopy groups  887

Lian, C. and R. Moschetti ℤ/rℤ-equivariant covers of ℙ^{1} with moving ramification

487

Lissitsin, A. See S. Ciaci Lizama, C. See L. Abadias Lü, R. See Y. Xue Melcher, R. See S. Geschke Mesyan, Z. See R. Hazrat Monti, V. See M. Avitabile Moon, J. See J. Brendel Morigi, M. See E. Detomi Moschetti, R. See C. Lian Müller, S. See Y. Hayut Newman, M. F. See M. Avitabile O’Brien, E. A. See M. Avitabile Pak, I. See P. Brändén Pan, J. See P. Li Papageorgiou, N. S. and C. Zhang Multiple ground-state solutions

with sign information for double-phase robin problems

419

Pitz, M. See S. Geschke Riedl, E. See I. Coskun Sah, A., M. Sawhney and Y. Zhao The cylindrical width of tran-

sitive sets

647

Sawhney, M. See A. Sah Šemrl, P. See M. Brešar Shumyatsky, P. See E. Detomi

Ueyama, K. Derived categories of skew quadric hypersurfaces  205

Veprev, G. Non-existence of a universal zero entropy system for nonperiodic amenable group actions

715

Vestfrid, I. A. ε-isometries between the positive cones of continuous functions spaces

989

Wibmer, M. Subgroups of free proalgebraic groups and Matzat’s conjecture for function fields

841

Wu, E. See J. D. Christensen Wu, J. See P. Li Xue, Y. and R. Lü Classification of simple bounded weight modules

of the Lie algebra of vector fields on ℂ^{n}

445

Zerbes, S. L. See A. Besser Zhang, C. See N. S. Papageorgiou Zhao, Y. See A. Sah